0. A sample of carbon dioxide, CO2, occupies 0.300L at 10 'c and 750 torr. What volume will the gas have at 30 'c and 750 torr?
1
Expert's answer
2020-05-24T02:59:49-0400

Here the pressure remains constant zo we have to calculate only the change of volume with respect to temperature.

From Charl's Law

V1/ T1 = V2 / T2

Here

V1 = 0.300 L

T1= 10°C

= 283 K

T2 = 30° C

= 303 K

So final volume, V2 will be

V2 = (0.300 × 303)/ 283

= 0.32 L